Circular No. 9126

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 E3 (WISE) An apparently asteroidal object found on WISE satellite images (dis-

covery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to be diffuse with a faint tail in p.a. ∼ 315^{o} on CCD images taken with the Magdalena Ridge Observatory 2.4-m f/8.9 reflector by W. H. Ryan and E. V. Ryan on Mar. 13.3 and 14.1 UT, measuring the object’s red magnitude to be 20.3 ± 0.4. Upon inquiry by the Central Bureau, A. Mainzer (Jet Propulsion Laboratory) notes notes an asymmetric coma of size ∼ 25′′ in 12- and 22-μm stacked WISE images.

2010 UT α_{2000} δ_{2000} Observer
Mar. 5.09910 4^{h}42^{m}43.^{s}55 +66^{o}52′38′′.1 WISE

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-F01.

T = 2010 Apr. 4.326 TT  ω = 49.^{o}974
Ω = 117.328 }2000.0
q = 2.27420 AU  i = 96.480

COMET P/2010 E4 (LINEAR) An apparently asteroidal object discovered in 2002 by the LINEAR

survey (discovery observation tabulated below), and designated 2002 CF_{140} on MPS 50320, has been found by R. A. Kowalski to show cometary appearance on Mount Lemmon 1.5-m reflector CCD images (identified at the Minor Planet Center; first observation tabulated below) taken in poor seeing during Mar. 15.40–15.46 UT. Four 75-s exposures show a slightly condensed round coma ∼ 15′′ in size, slightly elongated toward p.a. 275^{o}, which is the direction of a broad, straight tail 30′′ long.

Date UT α_{2000} δ_{2000} Mag. Observer
2002 Feb. 8.31900 9^{h}46^{m}37.^{s}48 +2^{o}53′48′′.1 18.6 LINEAR
2010 Mar. 15.40060 13 46 43.68 −5 53 21.0 19.0 Kowalski

The orbit on MPO 30668 requires a correction of ΔT = −0.8 day. The new astrometry (including 2010 Mar. 11 observations from La Sagra), new elliptical orbital elements by B. G. Marsden (epoch 2010 Jan. 4.0 TT, T = 2009 Dec. 23.4780 TT, q = 2.861297 AU, e = 0.250940, ω = 358^{o}.3346, Ω = 179^{o}.7271, i = 11^{o}.5150, equinox 2000.0, P = 7.47 years), and an ephemeris appear on MPEC 2010-F02.

2010 March 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green